P=$250-$0.15Q, TC=$25000+$10Q(The TC function does not include the firm's cost of capital)
a. In a unregulated enviroment, what price would this firm change, what output would be produced, what would total profits be, and what rate of return would the firm earn on its asset base?
1
Expert's answer
2017-08-15T08:40:07-0400
P=$250-$0.15Q, TC=$25000+$10Q
a. The price and quantity are optimal, when MR = MC = P, so:
MC = TC' = 10,
MR = TR' = (P*Q)' = 250 - 0.3Q,
250 - 0.3Q = 10,
0.3Q = 240,
Q = 800 units.
P = 250 - 0.15*800 = $130.
Total profit is:
TP = TR - TC = 130*800 - (25000 + 10*800) = $71000.